Shumarra exercises four dogs. Bing, Muskie, Cowboy,and Teddy. Teddy is 5 pounds heavier than Muskie. Cowboy is 8 pounds lighter than Bing, who is twice as heavy as Muskie. If Teddy weighs 25 pounds, what does each of the other dogs weigh?
Answer by jainenderkapoor(61)


Teddy = Muskie + 5
so Muskie = Teddy - 5 = 25 - 5 = 20 pounds
Bing = 2* Muskie = 2*20 = 40 pounds
Cowboy is 8 pounds lighter than Bing = 40 - 8 = 32 pounds



Hope the solution is clear to you. If you have any doubt you can contact me at jainenderkapoor@gmail.com I am also available for Online Tutoring at very reasonable rates.